OpenStudy (anonymous):

i need help with a polynomial word problem Suzie’s flower shop estimates that supplies costs for one month will be given by the polynomial 5x2 + x – 4. She estimates that utilities costs for one month is given by 8x – 7. Find and simplify an expression that would estimate Suzie’s total costs for one month.

OpenStudy (anonymous):

total cost = utilities cost + supplies cost = 5x^2 + x - 4 +8x -7 = 5x^2 +9x -11

OpenStudy (anonymous):

why do u add it and not multiply it?

OpenStudy (anonymous):

because it's total multiply if the function are same, 2+2+2+2 = 2*4 --> just for the same number/function

OpenStudy (anonymous):

o ok i get it! thanks!!

OpenStudy (anonymous):

not a problem :)